# Strassen $2\times 2$ Matrix Multiplication from a 3-dimensional Volume Form

### v1

@@ 1 Introduction
Strassen’s $2\times 2$ matrix multiplication algorithm [1] is a formula for multiplying $2\times 2$ matrices $a$ and $b$:
@@
This may be applied recursively to multiply $n\times n$ matrices in $O(n^{\log 7/\log 2})$ time, approximately $O(n^{2.81})$, opening a research field to which the book [2] provides an introduction. One line of research has focused on further improving this asymptotic complexity, notably [3], [4] achieving $O(n^{2.376})$ and several refinements, recently [5] and [6], approaching $O(n^{2.37})$. Another line of research has pursued decompositions of matrix multiplication tensors for other small matrix sizes such as $3\times 3$, $4\times 4$, etc. These have often involved numerical searches, such as the recent [7].
@@
Despite these advances, matrix multiplication algorithms faster than $O(n^{2.81})$ are “almost never implemented” [8], and practical evaluations such as [9] have continued favoring the Strassen $2\times 2$ algorithm. Known algorithms for other small matrix sizes struggle to significantly improve on it: the up-to-date table [10] shows complexity exponents clustering around 2.8. For $4\times 4$ matrix multiplication, the Strassen algorithm has tensor rank $497^{2}=$ 49, and that remained the state of the art for over 55 years until [7] and [11] lowered that from 49 to 48, achieving a complexity exponent of $2.79$. Moreover, known algorithms with substantially lower asymptotic complexity tend to have large constants in the $O$, as discussed in [12].
@@
The Strassen $2\times 2$ algorithm also stands out from a theoretical perspective: its tensor rank 7 is known to be optimal [13] and it is known to be essentially unique under that constraint [14]. By contrast, the tensor rank of $n\times n$ matrix multiplication is still unknown for all $n\geq 3$. For $n=3$, it is still only known to be between 19 and 23, see [15].
@@
Optimality and uniqueness make the Strassen $2\times 2$ algorithm a basic fact of 2-dimensional linear algebra. Such facts are expected to be simple and geometric. However, the original statement and proof of the Strassen algorithm are calculations on matrix coefficients. This has motivated a quest for geometric interpretations. The recent [16] and [8] in particular were inspirational to the present article, and [8] contains a survey of this endeavour, tracing it back to the years following the publication of the original Strassen article [1]. Other recent articles in this line of research include [17], [18], [19] and [20].
@@
The present article offers a geometric interpretation of the Strassen algorithm by addressing a more general question: is the Strassen algorithm an independent fact in multilinear algebra, or could it be related to a known fact? We derive it from the expansion of a 3-dimensional volume form into an antisymmetrized sum of $3!=6$ simple tensors. That expansion follows from the one-dimensionality of the space of antisymmetric $n$-forms, which is an abstract version of Cavalieri’s principle, the idea that the volume of a solid is unchanged by sliding parallel slices. As to the question of why specifically $2\times 2$ matrices, the answer is that as matrix multiplication is a tensor of order 3 on matrix spaces, interpreting it as a volume form requires a 3-dimensional matrix space, and the specific case of $2\times 2$ matrices gives us such a 3-dimensional matrix space by taking the quotient by multiples of the identity matrix: $13=2^{2}-$ 1.
@@
Acknowledgements. The author would like to thank Paolo d’Alberto and Zach Garvey at AMD for helpful comments.
@@ 2 Overview
Using the natural isomorphism $\mathrm{L}(V)^{*\otimes 3}\simeq\mathrm{L}(V^{\otimes 3})$, view the trilinear forms $\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$, $\mathrm{tr}(a_{1}a_{2}a_{3})$ and $\mathrm{tr}(a_{3}a_{2}a_{1})$ as respectively the permutations $\mathrm{id}$, (123) and (321) permuting the terms in $V^{\otimes 3}$ (Lemma 8). This allows viewing $h$ as the composition of $g$ with a linear map induced by the permutation (321) (Lemma 16), which allows transporting certain rank 6 decompositions of $g$ into rank 6 decompositions of $h$ (Proposition 18, our main result), yielding (Corollary 20)
@@ 3 Terminology and lemmas in tensor algebra
$\mathrm{L}(U_{1},W_{1})\otimes...\otimes\mathrm{L}(U_{n},W_{n})\simeq\mathrm{L}(U_{1}\otimes...\otimes U_{n},W_{1}\otimes...\otimes W_{n}).$
@@
$\iota\;$ $:\;U\otimes U^{*}\rightarrow\mathrm{L}(U),$ $v\otimes\lambda$ $\mapsto\iota(v\otimes\lambda)=(u\mapsto\lambda(u)v)$
$\iota^{*}\;$ $:\;U\otimes U^{*}\rightarrow\mathrm{L}(U)^{*},$ $v\otimes\lambda$ $\mapsto\iota^{*}(v\otimes\lambda)=(a\mapsto\lambda(a(v)))$
*\; $:\;\mathrm{L}(U)\rightarrow\mathrm{L}(U)^{*},$ $a$ $\mapsto a^{*}=(b\mapsto\mathrm{tr}(ab)).$
@@
[…]
@@
$=\iota(v\otimes\lambda)^{*}(\iota(u\otimes\mu))$ by Definition 1. $\;\qed$
@@
$(ab)^{*}(c)=\mathrm{tr}(abc)=\mathrm{tr}(bca)=\mathrm{tr}(bR_{a}(c))=b^{*}(R_{a}(c))=(b^{*}\circ R_{a})(c).\qed$
@@
The following lemma is classical and is encountered around Weyl invariant tensor theory ([21], [22]), which, among other things, establishes that the $t_{\sigma}$ span the space of $\mathrm{GL}(U)$-invariant tensors in $\mathrm{L}(U^{\otimes 3})$. We will not need any of that theory, but it is still useful context.
@@
For any vector space $U$, the images of $t_{\mathrm{id}}$, $t_{(123)}$, $t_{(321)}$ under the map $t\mapsto t^{*}$ from Definition 1 are the following trilinear forms, given by their values at any $a_{1}\otimes a_{2}\otimes a_{3}$ in $\mathrm{L}(U)^{\otimes 3}$:
@@
$t_{(123)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{1}a_{2}a_{3}),$
$t_{(321)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{3}a_{2}a_{1}).$
@@
$L_{t_{\sigma}}\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)=\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i}).$
@@
We have
$L_{t_{\sigma}}\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)$ $=t_{\sigma}\cdot\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})$
$=\bigotimes_{i=1,2,3}\iota(t_{\sigma}(u_{i})\otimes\zeta_{i})$
$=\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i}).\;\qed$
@@
$\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\circ L_{t_{\sigma}}=\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)}).$
@@
$\left(\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\circ L_{t_{\sigma}}\right)\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)$
$=\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\left(\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i})\right)$ by Lemma 9
@@
$=\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)})\right)\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right).$
@@ 4 Main results
$h$ $=(t_{\mathrm{id}}-t_{(123)})^{*}$ by Equation (13)
$=((t_{(123)}-t_{(321)})\cdot t_{(321)})^{*}$
$=(t_{(123)}-t_{(321)})^{*}\circ L_{t_{(321)}}$ by Lemma 6
$=g\circ L_{t_{(321)}}$ by Equation (12).∎
@@
$h=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)}).$ (14)
@@
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(i)})\right)\circ L_{t_{(321)}},$
@@
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)}).$ (15)
@@
$h(a_{1},a_{2},a_{3})=-tr(a_{1}a_{2}a_{3})=-\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1}).$ (16)
@@
$h(a_{1},a_{2},a_{3})=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma(123)(i)}(v_{i})\lambda_{i}(v_{\sigma(i)}).$ (17)
Since $\lambda_{i}(v_{i})=0$, the product in Equation (17) vanishes whenever $\sigma$ has a fixed point or $\sigma(123)$ has a fixed point. Thus the only $\sigma$ contributing to the sum is $\sigma=(123)$. Thus, Equation (17) simplifies to
$h(a_{1},a_{2},a_{3})=\alpha\prod_{i=1,2,3}\lambda_{(321)(i)}(v_{i})\lambda_{i}(v_{(123)(i)}).$ (18)
@@
$\alpha=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\cdot\;\qed$
@@
Two tensors $p,q$ in $\mathrm{L}(V)^{*\otimes 3}$ related to each other in the same way as $g$ and $h$ are related by Lemma 16, namely $q=p\circ L_{t_{(321)}}$, may still fail to have the same tensor rank if their tensor decompositions involve linear terms in $\mathrm{L}(V)^{*}$ that are not of rank one.
@@
Consider the counterexample of $p=t_{(123)}^{*}$ and $q=t_{\mathrm{id}}^{*}$. The same argument as in the proof of Lemma 16 yields $q=p\circ L_{t_{(321)}}$. As noted in Lemma 8, for $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$, we have $p(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1}a_{2}a_{3})$ and $q(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$. Thus, as tensors in $\mathrm{L}(V)^{*\otimes 3}$, $q$ has rank one but $p$ does not. ∎

### v2

@@ 1 Introduction
Strassen’s $2\times 2$ matrix multiplication algorithm [Strassen] is a formula for multiplying $2\times 2$ matrices $a$ and $b$:
@@
This may be applied recursively to multiply $n\times n$ matrices in $O(n^{\log 7/\log 2})$ time, approximately $O(n^{2.81})$, opening a research field to which the book [AlgebraicComplexityTheory] provides an introduction. One line of research has focused on further improving this asymptotic complexity, notably [StrassenLaser], [CoppersmithWinograd] achieving $O(n^{2.376})$ and several refinements, recently [williams2023newboundsmatrixmultiplication] and [alman2024asymmetryyieldsfastermatrix], approaching $O(n^{2.37})$. Another line of research has pursued decompositions of matrix multiplication tensors for other small matrix sizes such as $3\times 3$, $4\times 4$, etc. These have often involved numerical searches, such as the recent [novikov2025alphaevolvecodingagentscientific].
@@
Despite these advances, algorithms faster than $O(n^{2.81})$ are “almost never implemented” [IkenmeyerLysikov], and practical evaluations such as [dalberto2023strassensmatrixmultiplicationalgorithm] have continued favoring the Strassen $2\times 2$ algorithm. Known algorithms for other small matrix sizes struggle to significantly improve on it: the up-to-date table [TableFastMatmul] shows complexity exponents clustering around 2.8. For $4\times 4$ matrix multiplication, the Strassen algorithm has tensor rank $7^{2}=49$, and that remained the state of the art for over 55 years until [novikov2025alphaevolvecodingagentscientific] and [dumas2025noncommutativealgorithmmultiplying4x4] lowered that from 49 to 48, achieving a complexity exponent of $2.79$. Moreover, known algorithms with substantially lower asymptotic complexity tend to have large constants in the $O$, as discussed in [alman2024improvingleadingconstantmatrix].
@@
The Strassen $2\times 2$ algorithm also stands out from a theoretical perspective: its tensor rank 7 is known to be optimal [WINOGRAD1971381] and it is known to be essentially unique under that constraint [DEGROOTE19781]. By contrast, the tensor rank of $n\times n$ matrix multiplication is still unknown for all $n\geq 3$. For $n=3$, it is still only known to be between 19 and 23, see [Blaeser03].
@@
Optimality and uniqueness make the Strassen $2\times 2$ algorithm a basic fact of 2-dimensional linear algebra. Such facts are expected to be simple and geometric. However, the original statement and proof of the Strassen algorithm are calculations on matrix coefficients. This has motivated a quest for geometric interpretations. The recent [ChiantiniIkenmeyerLandsbergOttaviani] and [IkenmeyerLysikov] in particular were inspirational to the present article, and [IkenmeyerLysikov] contains a survey of this endeavour, tracing it back to the years following the publication of the original Strassen article [Strassen]. Other recent articles in this line of research include [ikenmeyer2025strassensalgorithmorbitflip], [Burichenko], [GrochowMooreGroupOrbits] and [GrochowMoore].
@@
The present article offers a geometric interpretation of the Strassen algorithm by addressing a more general question: is the Strassen algorithm an independent fact in multilinear algebra, or could it be related to a known fact? We derive it from the expansion of a 3-dimensional volume form into an antisymmetrized sum of $3!=6$ simple tensors. That expansion follows from the one-dimensionality of the space of antisymmetric $n$-forms, which is an abstract version of Cavalieri’s principle, the idea that the volume of a solid is unchanged by sliding parallel slices. As to the question of why specifically $2\times 2$ matrices, the answer is that as matrix multiplication is a tensor of order 3 on matrix spaces, interpreting it as a volume form requires a 3-dimensional matrix space, and the specific case of $2\times 2$ matrices gives us such a 3-dimensional matrix space by taking the quotient by multiples of the identity matrix: $3=2^{2}-1$.
@@
Acknowledgements. The author would like to thank Eugene Ha, Paolo d’Alberto and Zach Garvey for helpful comments.
@@ 2 Overview
Using the natural isomorphism $\mathrm{L}(V)^{*\otimes 3}\simeq\mathrm{L}(V^{\otimes 3})$, view the trilinear forms $\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$, $\mathrm{tr}(a_{1}a_{2}a_{3})$ and $\mathrm{tr}(a_{3}a_{2}a_{1})$ as respectively the permutations $\mathrm{id}$, (1 2 3) and (3 2 1) permuting the terms in $V^{\otimes 3}$ (Lemma 8). This allows viewing $h$ as the composition of $g$ with a linear map induced by the permutation (3 2 1) (Lemma 16), which allows transporting certain rank 6 decompositions of $g$ into rank 6 decompositions of $h$ (Proposition 18, our main result), yielding (Corollary 20)
@@ 3 Terminology and lemmas in tensor algebra
$\mathrm{L}(U_{1},W_{1})\otimes\ldots\otimes\mathrm{L}(U_{n},W_{n})\simeq\mathrm{L}(U_{1}\otimes\ldots\otimes U_{n},W_{1}\otimes\ldots\otimes W_{n}).$
@@
$\iota\;$ $\colon\;U\otimes U^{*}\rightarrow\mathrm{L}(U),$ $v\otimes\lambda$ $\mapsto\iota(v\otimes\lambda)=(u\mapsto\lambda(u)v)$
$\iota^{*}\;$ $\colon\;U\otimes U^{*}\rightarrow\mathrm{L}(U)^{*},$ $v\otimes\lambda$ $\mapsto\iota^{*}(v\otimes\lambda)=(a\mapsto\lambda(a(v)))$
$*\;$ $\colon\;\mathrm{L}(U)\rightarrow\mathrm{L}(U)^{*},$ $a$ $\mapsto a^{*}=(b\mapsto\mathrm{tr}(ab)).$
@@
$U\otimes U^{*}$$\mathrm{L}(U)$$\mathrm{L}(U)^{*}$$\scriptstyle{\iota}$$\scriptstyle{\iota^{*}}\scriptstyle{*}$ (8)
@@
$=\iota(v\otimes\lambda)^{*}(\iota(u\otimes\mu))$ by Definition 1.
@@
$(ab)^{*}(c)=\mathrm{tr}(abc)=\mathrm{tr}(bca)=\mathrm{tr}(bR_{a}(c))=b^{*}(R_{a}(c))=(b^{*}\circ R_{a})(c).$
@@
The following lemma is classical ([Markl2008], [Weyl]).
@@
For any vector space $U$, the images of $t_{\mathrm{id}}$, $t_{(1\,2\,3)}$, $t_{(3\,2\,1)}$ under the map $t\mapsto t^{*}$ from Definition 1 are the following trilinear forms, given by their values at any $a_{1}\otimes a_{2}\otimes a_{3}$ in $\mathrm{L}(U)^{\otimes 3}$:
@@
$t_{(1\,2\,3)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{1}a_{2}a_{3}),$
$t_{(3\,2\,1)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{3}a_{2}a_{1}).$
@@
$L_{t_{\sigma}}\Biggl(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\Biggr)=\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i}).$
@@
Both sides are equal to $t_{\sigma}\cdot\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i}).$
@@
$\Biggl(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\Biggr)\circ L_{t_{\sigma}}=\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)}).$
@@
$\Biggl(\Biggl(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\Biggr)\circ L_{t_{\sigma}}\Biggr)\Biggl(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\Biggr)$
$=\Biggl(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\Biggr)\Biggl(\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i})\Biggr)$ by Lemma 9
@@
$=\Biggl(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)})\Biggr)\Biggl(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\Biggr).$
@@ 4 Main results
$h$ $=(t_{\mathrm{id}}-t_{(1\,2\,3)})^{*}$ by Equation (13)
$=((t_{(1\,2\,3)}-t_{(3\,2\,1)})\cdot t_{(3\,2\,1)})^{*}$
$=(t_{(1\,2\,3)}-t_{(3\,2\,1)})^{*}\circ L_{t_{(3\,2\,1)}}$ by Lemma 6
$=g\circ L_{t_{(3\,2\,1)}}$ by Equation (12).∎
@@
$h=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma\cdot(1\,2\,3)(i)}).$ (14)
@@
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\Biggl(\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(i)})\Biggr)\circ L_{t_{(3\,2\,1)}},$
@@
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma\cdot(1\,2\,3)(i)}).$ (15)
@@
$h(a_{1},a_{2},a_{3})=-\mathrm{tr}(a_{1}a_{2}a_{3})=-\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1}).$ (16)
@@
$h(a_{1},a_{2},a_{3})=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma\cdot(1\,2\,3)(i)}(v_{i})\lambda_{i}(v_{\sigma(i)}).$ (17)
Since $\lambda_{i}(v_{i})=0$, the product in Equation (17) vanishes whenever $\sigma$ has a fixed point or $\sigma\cdot(1\,2\,3)$ has a fixed point. Thus the only $\sigma$ contributing to the sum is $\sigma=(1\,2\,3)$. Thus, Equation (17) simplifies to
$h(a_{1},a_{2},a_{3})=\alpha\prod_{i=1,2,3}\lambda_{(3\,2\,1)(i)}(v_{i})\lambda_{i}(v_{(1\,2\,3)(i)}).$ (18)
@@
$\alpha=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\cdot$
@@
Two tensors $p,q$ in $\mathrm{L}(V)^{*\otimes 3}$ related to each other in the same way as $g$ and $h$ are related by Lemma 16, namely $q=p\circ L_{t_{(3\,2\,1)}}$, may still fail to have the same tensor rank if their tensor decompositions involve linear terms in $\mathrm{L}(V)^{*}$ that are not of rank one.
@@
Consider the counterexample of $p=t_{(1\,2\,3)}^{*}$ and $q=t_{\mathrm{id}}^{*}$. The same argument as in the proof of Lemma 16 yields $q=p\circ L_{t_{(3\,2\,1)}}$. As noted in Lemma 8, for $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$, we have $p(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1}a_{2}a_{3})$ and $q(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$. Thus, as tensors in $\mathrm{L}(V)^{*\otimes 3}$, $q$ has rank one but $p$ does not. ∎